# A Quadratically-Constrained Convex Approximation for the AC Optimal Power Flow

### v1

@@ III-B1 Optimality gap
$\boldsymbol{\bar{p}}^{\rm CVX}\in$ $\operatorname*{arg\,min}_{\Xi}\quad\lVert\hat{\boldsymbol{p}}^{\rm CVX}-\boldsymbol{p}^{g}\rVert_{2}^{2}$ (5a)
@@
where the superscript $\rm CVX$ represents the corresponding linear/convex approximation/relaxations, i.e., $\rm CVX=\{QCAC,TS,SOC,$ $\rm SDP\}$, and $\boldsymbol{\bar{p}}^{\rm CVX}$ denotes the Euclidean projection of $\hat{\boldsymbol{p}}^{\rm CVX}$ onto the constraint set defined by (1b)-(1n). The above optimization problem projects the active power generation dispatch to the feasible set of the AC-OPF problem. Note that Problem (5) is a nonconvex problem, and we solve it using Ipopt, which is a local solver that produces feasible solutions but not necessarily the closest feasible projection.
The optimality gap is determined as follows:
@@ III-B2 Distance to feasibility
[…]
III-B3 Runtime
For the four models, we report the time to solve the relaxations/approximations in addition to the time to find the closest feasible dispatch, i.e., time to solve (5).
III-C Results for small- and medium-scale instances
Table I presents the results for the small to medium-scale instances using the solution of the base case as a convexification point. The SDP relaxation is solved for instances up to 118 buses since the solution of larger instances violates the time limit of 600 seconds. The results show that the proposed approximation is competitive in terms of the optimality gap with respect to the other approximate models, particularly in the larger instances. In terms of distance to feasibility, the QCAC model notably outperforms the other three models in the majority of test cases, showing its ability to render a more secure generation dispatch. Notice that the quality of the solutions of the proposed approximation might (and often does) depend on how far the operating point is with respect to the convexification point.
III-D Results for large-scale instances
@@
Table II summarizes the correlation coefficients and mean/maximum errors of voltage magnitudes, phase angles, and active/reactive power generation. The QCAC model shows high accuracy across the three systems, showing its ability to properly convexify the nonlinearities of the AC-OPF problem. The reactive power is the most challenging variable in terms of both correlation coefficients and absolute errors. Figure 2 illustrates the high correlation of the four variables for an instance of the largest test system.
TABLE II: Correlation Coefficients, and Mean and Maximum Errors for Selected PGLIB Large-Scale Test Cases
@@
Figure 4 shows the time performance profile for the three systems for the QCAC model and the nonconvex AC-OPF problem. The proposed model outperforms the nonconvex model in terms of solution time and reduced runtime variability. The largest system shows the highest runtime speedup, which empirically demonstrates the scalability of the proposed convexified model.
@@
IV Case studies
In this section, we illustrate the application of the proposed approximation in two case studies: optimal reactive power flow with discrete controllers and a PV hosting capacity problem. The case studies indicate that the proposed QCAC model provides an accurate convex approximation of the power flow equations, which is particularly useful when other approximations/relaxations are slow, inaccurate, or neglect voltage- and reactive-power-related variables and constraints. The numerical experiments were implemented considering the same setup as in Section III.
IV-A Optimal Reactive Power Flow with Discrete Controllers
The optimal reactive power flow (ORPF) problem, a mixed-integer nonconvex problem that derives from the OPF problem, aims at minimizing the generation costs by adjusting the active and reactive power generation setpoints as well as by operating discrete controllers, e.g., switchable shunt reactive power compensation [32, 33]. Operating such discrete controllers aims at acting on the nodal voltages by adjusting the reactive power injections to reduce the generating costs, requiring a detailed representation of both variables, i.e., voltage magnitudes and reactive power flows. Such requirement hinders the use of linear approximations of the power flow equations, such as the DC power flow, that neglect the aforementioned variables.
@@ IV-A Optimal Reactive Power Flow with Discrete Controllers
The nonconvexities of Problem (6) arise from (i) the bilinear equalities, (6c), (ii) the bilinear and quadratic equalities, (1l)-(1n), and (iii) the discrete variables representing the steps of shunt compensation (6e). The first sources of nonconvexities can be exactly linearized since $b_{i}^{\rm sh}$ are discrete, and both variables in the bilinear terms are bounded [34, § 9.2]. The interaction of binary variables with the aforementioned nonconvexities makes impractical the solution of the original model for medium or large-scale systems.
@@
To find the convexification point, $V_{i}^{\rm re}$ and $V_{i}^{\rm im}$, we solve the continuous relaxation of Problem (6). The experimental results study the benefits and accuracy of the convexified problem with respect to the original mixed-integer nonconvex formulation and the one using a SOC relaxation of the power flow equations. We use a 30-bus test system [9] under a normal operating condition (TYP) to test the performance of the QCAC approximation.
@@
Table III presents the numerical results for the ORPD problem. Gurobi was unable to find a feasible solution to the original MINLP model within 2 hours. Hence, we present the solution of its continuous relaxation, which provides a valid lower bound on the cost. To compute the cost and savings of the QCAC and SOC models, we fix the discrete variables obtained by the corresponding mixed-integer convex approximations and solve the restricted Problem (6), which provides a feasible solution. Note that the nonconvex problems are solved with Knitro, which cannot guarantee global optimality.
The proposed approximation outperforms the one obtained by using the MISOCP relaxation in terms of cost savings, reducing the production cost by 19.43% compared to the 14.49% of the SOC relaxation. Note as well that techniques such as Benders decomposition, outer approximation, and an active set strategy can be used to reduce the computational burden of solving the problem and enable the scalability [35, 36] of the SOC- and QCAC-based ORPD problems.
TABLE III: Results Summary: Proof-of-concept ORPD Problem
@@
IV-B PV Hosting Capacity Problem
We illustrate the application of the QCAC approximation to a 533-node radial distribution network whose data can be retrieved from Matpower [37]. Particularly, we focus on an application where the SOC relaxation of the power flow equations renders highly inaccurate solutions [38]. One such application is the PV hosting capacity problem, which aims to determine the maximum total PV generation that can be accommodated by a distribution feeder without violating its operation limits [39]. The PV hosting capacity problem is formulated as follows:
@@ IV-B PV Hosting Capacity Problem
$\;\sum_{k\in\mathcal{G}_{i}}p_{k}^{\rm g}+p_{i}^{\rm pv}-P_{i}^{\rm d}-G_{i}^{\rm sh}c_{ii}=\smashoperator[]{\sum_{(i,j)\in\mathcal{E}_{i}}^{}}p_{ij}+\smashoperator[]{\sum_{(j,i)\in\mathcal{E}_{i}}^{}}p_{ji},$
$\hskip 184.9429pt\forall i\in\mathcal{N},$ (7b)
@@
Table IV summarizes the results of the three models, i.e., nonconvex, SOC, and QCAC. For all the penalty coefficients, the proposed model offers a speedup of 33$\approx$ 33$\times$ 33 × on average with the best cost error of $\approx 1\%$ whereas the SOC relaxation shows a higher speed up of 45×45$\times$45 ×, but with a significantly higher cost error of $\approx 245\%$.
TABLE IV: Result Summary: PV hosting capacity problem
@@
V Conclusion
In this paper, we propose a quadratically-constrained convex approximation for the AC-OPF problem that relies solely on a convexification point, without the need for additional assumptions. The approach is based on a difference-of-convex reformulation of the power flow equations, followed by a convexification procedure using a first-order Taylor series expansion of the resulting concave terms. Our extensive numerical results demonstrate the advantages of this model, including a reduced optimality gap, near-feasible solutions, and smaller solution errors compared to the nonconvex AC formulation. Furthermore, two proof-of-concept case studies illustrate the potential benefits of the proposed approximation over the second-order conic and semidefinite relaxations of the power flow equations.

### v2

@@ 3.2 Metrics
where the superscript $\rm CVX$ represents the corresponding linear/convex approximation/relaxations, i.e., $\rm CVX=\{QCAC,TS,SOC,$ $\rm SDP\}$, and $\boldsymbol{\bar{p}}^{\rm CVX}$ denotes the Euclidean projection of $\hat{\boldsymbol{p}}^{\rm CVX}$ onto the constraint set defined by (1b)-(1n).
The above optimization problem projects the active power generation dispatch to the feasible set of the AC-OPF problem. Note that Problem (5) is a nonconvex problem, and we solve it using Ipopt, which is a local solver that produces feasible solutions but not necessarily the closest feasible projection. The optimality gap is determined as follows:
@@
$\text{Distance to feasibility}\,(\textrm{p.u.})=\sqrt{\frac{1}{\lvert\mathcal{G}\rvert}\sum_{k\in\mathcal{G}}\left(\overline{p}_{k}^{\rm CVX}-\hat{p}_{k}^{\rm CVX}\right)^{2}}.$
Runtime
For the four models, we report the time to solve the relaxations and approximations in addition to the time to find the closest feasible dispatch, i.e., time to solve (5).
3.3 Results for small- and medium-scale instances
Table 1 presents the results for the small to medium-scale instances using the solution of the base case as a convexification point. The SDP relaxation is solved for instances up to 118 buses since the solution of larger instances violates the time limit of 600 seconds. The results show that the proposed approximation is competitive in terms of the optimality gap with respect to the other approximate models, particularly in the larger instances. In terms of distance to feasibility, the QCAC model notably outperforms the other three models in the majority of test cases, showing its ability to render a more secure generation dispatch. Notice that the quality of the solutions of the proposed approximation might (and often does) depend on how far the operating point is with respect to the convexification point.
Table 1: Optimality Gaps, Distances to Feasibility, and Runtimes for Selected PGLIB Test Cases
Optimality Gap (%) Distance to feasibility (p.u.) Runtime (s) Test case QCAC TS SOC SDP QCAC TS SOC SDP QCAC TS SOC SDP pglib_opf_case30 0.06256 0.27451 0.66478 0.00680 0.14951 0.39028 0.29270 0.00004 0.03 0.03 0.04 0.82 pglib_opf_case39 0.27498 2.60725 3.26863 0.21014 0.01517 0.80445 1.09835 0.07128 0.03 0.03 0.04 2.82 pglib_opf_case57 0.01706 0.05375 0.00402 0.00092 0.01557 0.61258 0.01900 0.01852 0.05 0.04 0.04 12.39 pglib_opf_case89 1.49676 7.79117 1.14178 7.15318 0.36341 1.29853 0.52210 1.18675 0.37 0.35 0.51 155.7 pglib_opf_case118 0.37494 10.01670 2.89311 2.06640 0.00300 0.33364 0.09849 0.06768 0.18 0.17 0.24 638.3 pglib_opf_case179 1.21317 0.04374 2.46625 - 0.02196 0.004736 1.32447 - 0.58 0.36 0.19 - pglib_opf_case300 1.53902 6.82574 1.23162 - 0.18725 0.39919 0.48233 - 1.83 2.21 2.16 - pglib_opf_case500 0.31590 2.06582 0.13161 - 0.01643 0.14463 0.01076 - 1.05 1.00 1.39 - pglib_opf_case793 0.81594 21.87617 3.85507 - 0.00300 0.16165 0.09789 - 4.81 4.01 4.32 - pglib_opf_case1354 1.35456 57.87368 4.12804 - 0.01238 1.61108 0.54144 - 4.32 4.32 5.09 - The highlighted values correspond to the best out of the QCAC, TS, SOC, and SDP models.
3.4 Results for large-scale instances
@@
Table 2 summarizes the correlation coefficients and mean/maximum errors of voltage magnitudes, phase angles, and active/reactive power generation. The QCAC model shows high accuracy across the three systems, showing its ability to properly convexify the nonlinearities of the AC-OPF problem. The reactive power is the most challenging variable in terms of both correlation coefficients and absolute errors. Figure 2 illustrates the high correlation of the four variables for an instance of the largest test system.
Table 2: Correlation Coefficients, and Mean and Maximum Errors for Selected PGLIB Large-Scale Test Cases
@@
Figure 4 shows the time performance profile for the three systems for the QCAC model and the AC-OPF problem. The proposed model outperforms the nonconvex model in terms of solution time and reduced runtime variability. The largest system shows the highest runtime speedup, which empirically demonstrates the scalability of the proposed convexified model.
@@
4 Case study
In this section, we illustrate the application of the proposed approximation in the optimal reactive power flow with discrete controllers. The case study indicate that the proposed QCAC model provides an accurate convex approximation of the power flow equations, which is particularly useful when other approximations/relaxations are slow, inaccurate, or neglect voltage- and reactive-power-related variables and constraints. The numerical experiments were implemented considering the same setup as in Section 3.
The optimal reactive power flow (ORPF) problem, a mixed-integer nonconvex problem that derives from the OPF problem, aims at minimizing the generation costs by adjusting the active and reactive power generation setpoints as well as by operating discrete controllers, e.g., switchable shunt reactive power compensation Yang2017 ; Constante2021 . Operating such discrete controllers aims at acting on the nodal voltages by adjusting the reactive power injections to reduce the generating costs, requiring a detailed representation of both variables, i.e., voltage magnitudes and reactive power flows. Such requirement hinders the use of linear approximations of the power flow equations, such as the DC power flow, that neglect the aforementioned variables.
@@ 4 Case study
The nonconvexities of Problem (6) arise from (i) the bilinear equalities, (6c), (ii) the bilinear and quadratic equalities, (1l)-(1n), and (iii) the discrete variables representing the steps of shunt compensation (6e). The first sources of nonconvexities can be exactly linearized since $b_{i}^{\rm sh}$ are discrete, and both variables in the bilinear terms are bounded (Williams2013, , § 9.2). The interaction of binary variables with the aforementioned nonconvexities makes impractical the solution of the original model for medium or large-scale systems.
@@
To find the convexification point, $V_{i}^{\rm re}$ and $V_{i}^{\rm im}$, we solve the continuous relaxation of Problem (6). The experimental results study the benefits and accuracy of the convexified problem with respect to the original mixed-integer nonconvex formulation and the one using a SOC relaxation of the power flow equations. We use a 30-bus test system Alsac1974 under a normal operating condition (TYP) to test the performance of the QCAC approximation.
@@
Table 3 presents the numerical results for the ORPD problem. Gurobi was unable to find a feasible solution to the original MINLP model within 2 hours. Hence, we present the solution of its continuous relaxation, which provides a valid lower bound on the cost. To compute the cost and savings of the QCAC and SOC models, we fix the discrete variables obtained by the corresponding mixed-integer convex approximations and solve the restricted Problem (6), which provides a feasible solution. Note that the nonconvex problems are solved with Knitro, which cannot guarantee global optimality.
The proposed approximation outperforms the one obtained by using the MISOCP relaxation in terms of cost savings, reducing the production cost by 19.43% compared to the 14.49% of the SOC relaxation. Note as well that techniques such as Benders decomposition, outer approximation, and an active set strategy can be used to reduce the computational burden of solving the problem and enable the scalability Constante2024 ; Jabr2024 of the SOC- and QCAC-based ORPD problems.
Table 3: Results Summary: Proof-of-concept ORPD Problem
@@
4.1 Photovoltaic Hosting Capacity Problem
We illustrate the application of the QCAC approximation to a 533-node radial distribution network whose data can be retrieved from Matpower Zimmerman2011 . Particularly, we focus on an application where the SOC relaxation of the power flow equations renders highly inaccurate solutions Low2014 . One such application is the photovoltaic (PV) hosting capacity problem, which aims to determine the maximum total PV generation that can be accommodated by a distribution feeder without violating its operation limits Jha2022 . The PV hosting capacity problem is formulated as follows:
@@ 4.1 Photovoltaic Hosting Capacity Problem
$\;\sum_{k\in\mathcal{G}_{i}}p_{k}^{\rm g}+p_{i}^{\rm pv}-P_{i}^{\rm d}-G_{i}^{\rm sh}c_{ii}=\smashoperator[]{\sum_{(i,j)\in\mathcal{E}_{i}}^{}}p_{ij}+\smashoperator[]{\sum_{(j,i)\in\mathcal{E}_{i}}^{}}p_{ji},\forall i\in\mathcal{N},$ (7b)
@@
Table 4 summarizes the results of the three models, i.e., nonconvex, SOC, and QCAC. For all the penalty coefficients, the proposed model offers a speedup of $\approx 33\times$ on average with the best cost error of $\approx 1\%$ whereas the SOC relaxation shows a higher speed up of $45\times$, but with a significantly higher cost error of $\approx 245\%$.
Table 4: Result Summary: PV hosting capacity problem
@@
5 Conclusion
In this paper, we propose a quadratically-constrained convex approximation for the AC-OPF problem that relies solely on a convexification point, without the need for additional assumptions. The approach is based on a difference-of-convex reformulation of the power flow equations, followed by a convexification procedure using a first-order Taylor series expansion of the resulting concave terms. Our extensive numerical results demonstrate the advantages of this model, including a reduced optimality gap, near-feasible solutions, and smaller solution errors compared to the AC formulation. Furthermore, a case study illustrates the potential benefits of the proposed approximation over the second-order conic relaxation of the power flow equations for applications beyond the optimal power flow problem.
@@ 5 Conclusion
Declarations
Funding This research was supported by the Davidson School of Chemical Engineering at Purdue University and the Office of Naval Research under grant no. N000142412641.
Conflict of interest/Competing interests The authors declare that they have no competing interests.
Ethics approval and consent to participate Not applicable.
Consent for publication Not applicable.
Data availability The data that support the findings of this study are available from the corresponding author upon reasonable request.
Materials availability Not applicable.
Code availability Not applicable.
Author contributions G.C.F. developed the proposed model, conducted the theoretical analysis, and implemented the numerical experiments. C.L. provided guidance on the methodology, supervised the research progress, and contributed to interpreting the results. G.C.F. and C.L. jointly wrote the manuscript and prepared all figures and tables. All authors reviewed and approved the final manuscript.